# MARIA: a Multimodal Transformer Model for Incomplete Healthcare Data

### v1

@@ Abstract
In healthcare, the integration of multimodal data is pivotal for developing comprehensive diagnostic and predictive models. However, managing missing data remains a significant challenge in real-world applications. We introduce MARIA (Multimodal Attention Resilient to Incomplete datA), a novel transformer-based deep learning model designed to address these challenges through an intermediate fusion strategy. Unlike conventional approaches that depend on imputation, MARIA utilizes a masked self-attention mechanism, which processes only the available data without generating synthetic values. This approach enables it to effectively handle incomplete datasets, enhancing robustness and minimizing biases introduced by imputation methods. We evaluated MARIA against state-of-the-art machine learning and deep learning models across diagnostic and prognostic tasks. The results demonstrate that MARIA outperforms existing methods in terms of performance and resilience to varying levels of data incompleteness, underscoring its potential for critical healthcare applications.
@@ 1 Introduction
The MARIA (Multimodal Attention Resilient to Incomplete datA) model introduced in this work is designed to address the challenges of incomplete multimodal data. By employing an intermediate fusion strategy, MARIA combines modality-specific encoders with a shared attention-based encoder to effectively manage missing data. Unlike traditional methods that depend on data imputation to fill in missing entries, MARIA focuses exclusively on the available features, utilizing a masked self-attention mechanism to process observed information without generating synthetic data. This approach enhances both robustness and accuracy, while reducing biases typically introduced by imputation techniques.
@@ 3.1 Model
MARIA is specifically designed to be resilient to incomplete data and modalities without relying on imputation. It employs intermediate fusion, using modality-specific encoders and a shared encoder with a masked self-attention mechanism to combine latent representations while effectively managing missing data. The architecture incorporates multiple modality-specific encoders for each data modality. In this work, we focus on multimodal problems where the modalities represent tabular data describing various aspects of a patient’s condition. Thus, MARIA utilizes separate NAIM [7] modules as modality-specific encoders (Figure 2). These encoders integrate a masked multi-head attention mechanism that selectively focuses on available features within each modality while ignoring missing ones. This mechanism ensures robustness by excluding absent features from attention computations.
@@
This architecture qualifies MARIA as an intermediate fusion model, performing fusion at the latent representation level. By dynamically optimizing all encoders, MARIA balances the contributions of different modalities and adapts to maximize their utility during training and inference. The use of a masked multi-head attention mechanism ensures the model focuses adaptively on informative parts of the input, ignoring missing data. This approach allows each modality to contribute based on its completeness, resulting in accurate and reliable multimodal representations.
@@ 4.1 Data
We evaluated MARIA and its competitor models on two publicly available datasets across eight diagnostic and prognostic tasks (details in Table 1). These datasets represent real-world scenarios where patient data is often incomplete, highlighting the need for methods that are resilient to missing information.
@@
The first dataset was obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu) [12]. The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging, positron emission tomography, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s disease (AD), with respect to the Cognitively Normal (CN) group. For our study, we used four baseline data modalities: Assessment (cognitive and neuropsychological scores), Biospecimen (CSF, ApoE genotyping, and lab data), Image analysis (MRI and PET neuroimaging biomarkers), and Subject Characteristics (family history, demographics). We analyzed data from ADNI 1, GO, 2, and 3 phases. Diagnostic tasks included binary classification (CN vs. AD) and ternary classification (CN vs. AD vs. MCI), reflecting real clinical differentiation scenarios. Additionally, prognostic tasks aimed to predict whether treatment intervention might be necessary at specific future points (12, 24, 36, and 48 months post-recruitment). These tasks classified patients into CN, MCI, and Dementia categories. Table 1 provides details on patient distributions for both baseline and follow-up classifications.
@@ 4.2 Competitors
We conduct an extensive comparison of our methodology against early, late, and intermediate fusion approaches that use missing data imputation as a preprocessing step before model training. Additionally, we benchmark against tree-based models that manage missing values using the Missing In Attributes (MIA) strategy. We choose not to include generative approaches, as these are primarily developed for imaging modalities rather than tabular data, and they introduce additional complexity and randomness, which can hinder reproducibility. Instead, we focus on interpretable, efficient approaches widely adopted in clinical settings, where model simplicity and reliability are crucial. Our analysis includes a total of distinct competitor models, each combined with the k-Nearest Neighbors (kNN) imputation technique using default parameters, as this method outperformed other imputation strategies in prior studies [5, 7, 14].
@@ Acknowledgment
This work was partially founded by: i) Università Campus Bio-Medico di Roma under the program “University Strategic Projects” within the project “AI-powered Digital Twin for next-generation lung cancEr cAre (IDEA)”; ii) from PRIN 2022 MUR 20228MZFAA-AIDA (CUP C53D23003620008); iii) from PRIN PNRR 2022 MUR P2022P3CXJ-PICTURE (CUP C53D23009280001); iv) from PNRR MUR project PE0000013-FAIR.

### v2

@@ Abstract
In healthcare, the integration of multimodal data is pivotal for developing comprehensive diagnostic and predictive models. However, managing missing data remains a significant challenge in real-world applications. We introduce MARIA (Multimodal Attention Resilient to Incomplete datA), a novel transformer-based deep learning model designed to address these challenges through an intermediate fusion strategy. Unlike conventional approaches that depend on imputation, MARIA utilizes a modified masked self-attention mechanism, which processes only the available data without generating synthetic values. This approach enables it to effectively handle incomplete datasets, enhancing robustness and minimizing biases introduced by imputation methods. We evaluated MARIA against state-of-the-art machine learning and deep learning models across diagnostic and prognostic tasks. The results demonstrate that MARIA outperforms existing methods in terms of performance and resilience to varying levels of data incompleteness, underscoring its potential for critical healthcare applications.
@@ 1 Introduction
The MARIA (Multimodal Attention Resilient to Incomplete datA) model introduced in this work is designed to address the challenges of incomplete multimodal data. By employing an intermediate fusion strategy, MARIA combines modality-specific encoders with a shared attention-based encoder to effectively manage missing data. Unlike traditional methods that rely on data imputation to fill in missing entries, MARIA focuses exclusively on the available features, utilizing a modified masked self-attention mechanism to process observed information only, while completely masking out the contribution of missing information. This approach enhances both robustness and accuracy, while reducing biases typically introduced by imputation techniques.
@@ 3.1 Model
MARIA is specifically designed to be resilient to incomplete data and modalities without relying on imputation. It employs intermediate fusion, using modality-specific encoders and a shared encoder with a masked self-attention mechanism to combine latent representations while effectively managing missing data. The architecture incorporates multiple modality-specific encoders for each data modality. In this work, we focus on multimodal problems where the modalities represent tabular data describing various aspects of a patient’s condition. Thus, MARIA utilizes separate NAIM [7] modules as modality-specific encoders (Figure 2). These encoders integrate a modified masked multi-head attention mechanism that selectively focuses on available features within each modality while completely ignoring missing ones. This mechanism ensures robustness by excluding absent features from attention computations.
@@
This architecture qualifies MARIA as an intermediate fusion model, performing fusion at the latent representation level. By dynamically optimizing all encoders, MARIA balances the contributions of different modalities and adapts to maximize their utility during training and inference. The use of a masked multi-head attention mechanism ensures the model focuses adaptively on informative parts of the input, completely ignoring missing data. This approach allows each modality to contribute based on its completeness, resulting in accurate and reliable multimodal representations.
@@ 4.1 Data
We evaluated MARIA and its competitor models on two publicly available datasets across eight diagnostic and prognostic tasks (details in Table 1). These tabular datasets represent real-world scenarios where patient data is often incomplete, highlighting the need for methods that are resilient to missing information.
@@
The first dataset was obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu) [12]. The ADNI was launched in 2003 as a public-private partnership, led by Principal Investigator Michael W. Weiner, MD. The primary goal of ADNI has been to test whether serial magnetic resonance imaging, positron emission tomography, other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s disease (AD), with respect to the Cognitively Normal (CN) group. For our study, we used tabular data derived from four baseline modalities: Assessment (cognitive and neuropsychological scores), Biospecimen (CSF, ApoE genotyping, and lab data), Image analysis (MRI and PET neuroimaging biomarkers), and Subject Characteristics (family history, demographics). We analyzed data from ADNI 1, GO, 2, and 3 phases. Diagnostic tasks included binary classification (CN vs. AD) and ternary classification (CN vs. AD vs. MCI), reflecting real clinical differentiation scenarios. Additionally, prognostic tasks aimed to predict whether treatment intervention might be necessary at specific future points (12, 24, 36, and 48 months post-recruitment). These tasks classified patients into CN, MCI, and Dementia categories. Table 1 provides details on patient distributions for both baseline and follow-up classifications.
@@ 4.2 Competitors
We conduct an extensive comparison of our methodology against early, late, and intermediate fusion approaches that use missing data imputation as a preprocessing step before model training. Additionally, we benchmark against tree-based models that manage missing values using the Missing In Attributes (MIA) strategy. We choose not to include generative approaches, as these are primarily developed for imaging modalities rather than tabular data, and they introduce additional complexity and randomness, which can hinder reproducibility. Instead, we focus on interpretable, efficient approaches widely adopted in clinical settings, where model simplicity and reliability are crucial. Our analysis includes a total of distinct competitor models, each combined with the k-Nearest Neighbors (kNN) imputation technique, as this method outperformed other imputation strategies in prior studies [5, 7, 14]. For all models, we used default hyperparameters. Indeed, as shown in [15], in line with the “No Free Lunch” theorem, the authors empirically observe that, in many cases, tuning hyperparameters does not lead to significantly better performance compared to the default values suggested in the literature.
@@ Acknowledgment
This work was partially founded by: i) Università Campus Bio-Medico di Roma under the program “University Strategic Projects” within the project “AI-powered Digital Twin for next-generation lung cancEr cAre (IDEA)”; ii) from PRIN 2022 MUR 20228MZFAA-AIDA (CUP C53D23003620008); iii) from PRIN PNRR 2022 MUR P2022P3CXJ-PICTURE (CUP C53D23009280001); iv) from PNRR MUR project PE0000013-FAIR; v) Project ECS 0000024 Rome Technopole (CUP C83C22000510001).